Thomas’ Calculus Early Transcendentals 12ed

Homework Assignment for Chapter 06

1. Section 6.1: problems 5, 55, 63.

2. Section 6.2: problem 39.

3. Section 6.3: problems 21, 25.

4. Section 6.4: problem 5.

5. Chap 6, Additional and Advanced Exercises: problems 4.